OpenStudy (anonymous):

The price of a Ipod dropped from $299.99 to $180.55. What was the percent decrease in price. (Rounded to nearest hundredth percent).

OpenStudy (anonymous):

$(299.99-180.55)=119.44 dollars Percent decrease= (119.44/299.99)*100=39.81 % (approx)

OpenStudy (anonymous):

OK thank you I just noticed different answers and 39.81 was what I was getting

OpenStudy (anonymous):

your welcome